OpenStudy (anonymous):

Thanks for helping! I simplified a pre calc problem into cos(x)=1/sqrt(2) From here I'm stuck, as I need to write it in terms of pi. I would really appreciate any help, thanks!

OpenStudy (anonymous):

@Hero @iambatman @Kainui @jim_thompson5910 @ganeshie8 @zepdrix @Nnesha

OpenStudy (anonymous):

Which value of x makes \(cos(x) = \frac{1}{\sqrt{2}} = \frac{\sqrt{2}}{2}\) ?

OpenStudy (anonymous):

remember the basic cos, sin and tan of 30, 45 and 60 deg?

Nnesha (nnesha):

you can't hve the radical sign at the bottom so multiply top and bottom both with square root 2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

oh, wait, I remember this now. Thanks for the help guys!

OpenStudy (anonymous):

No problem ^^